# references.bib

### v1

@@ III-E Physics
Physical plausibility of reconstructed objects and scenes is essential if any physics simulation is to be performed. For rigid bodies, physical plausibility entails aspects of the overall scene composition, such as object collision and stability, as well as physical properties of individual objects themselves.
@@
On the scene level, many approaches formulate loss terms that are incorporated into gradient-based optimisation. For instance, SDF-based representations allow querying whether a particular point in space belongs to the volume of more than one object. As the SDF representation itself is typically optimised with gradient descent, adding a penalty term for doubly occupied space into the overall optimisation is straightforward [wu_object-compositional_2022, zhang_holistic_2021]. [li_rico_2023] and [chatterjee_3d-scene-former_2024] also regularise collision between objects and scene boundaries and background walls by relying on SDFs and bounding boxes, respectively. [wu_clusteringsdf_2025] and [hassena_objectcarver_2024] take a slightly different approach by not discouraging existing collisions between objects, but instead clustering and decomposing a scene representation into individual objects which naturally addresses object overlap. [yao_cast_2025] propose a complex setup that first queries a vision-language model to produce a scene graph containing object-object relations, such as one object supporting another or objects generally being in close contact with one another. Based on these pairwise graph relations, the authors select loss terms from a set of handcrafted loss functions that are then used to optimise physical scene coherence via gradient descent. [huang_midi_2024] do not formulate explicit loss terms but instead resolve collisions by coupling the denoising process of multiple objects in their proposed 3D diffusion model, leading to fewer violations.
@@
Instead of handcrafted loss-terms, some approaches resort to differentiable physics simulation. This allows penalising anything that would result in undesirable behaviour during physics simulation, including object collision and object stability -- even involving multiple objects at once. [yan_phycage_2024] dissect single objects into several components, such as a frog wearing a scarf, and optimise component shapes to resolve collision and object stability. [ni_phyrecon_2024] perform a similar process on the scene level by reconstructing individual objects and adapting their position and shape such that the resulting scene is physically stable.
@@
Some works do not consider entire scenes or compositional objects but instead concentrate on individual objects only. [mezghanni_physically-aware_2021] learn a stability predictor that is used for guiding object reconstruction. [guo_physically_2024] formulate differentiable loss terms to optimise stability of single objects. [chen_atlas3d_2024] use differentiable simulation during text-to-3D generation for generating stable objects, and [mezghanni_physical_2022] augment an SDF decoder with a differentiable simulator for object stability.
@@
Beyond stability and collision, researchers have also looked at estimating object articulations [chen_urdformer_2024, le_articulate-anything_2024, li_dragapart_2025] and material properties [boss_sf3d_2024, feng_pie-nerf_2024, guo_physically_2024]. While certainly useful for digital twinning of more complex environments, these aspects go beyond the scope of our study of whether reconstructing physically coherent scenes in simulation for robotics is feasible with current open-vocabulary 3D reconstruction methods.
@@ III-F Training Data
Modern AI is heavily data-driven, relying on ever-larger datasets [schuhmann_laion-5b_2022, kirillov_segment_2023, patel_introduction_2020]. In contrast to text and image data, 3D data is not as prevalent on the internet which is the reason why 3D datasets have historically been smaller-scale than their text or image counterparts.
@@
Most existing large-scale 3D datasets are primarily object-centric rather than scene-centric. This focus on isolated objects fails to capture important contextual relationships between objects in realistic environments. For robotics applications seeking to create accurate digital twins of entire scenes, having training data that represents realistic object arrangements, support relationships, and partial occlusions is crucial. Scene-level data would allow reconstruction models to learn typical object configurations and physical constraints that govern real-world environments. Without such contextual learning, models struggle to produce physically plausible reconstructions of complex scenes, leading to the stability and collision issues we observe in our evaluations.
Object Datasets
@@
ShapeNet [chang_shapenet_2015] has been very influential ever since its introduction, offering around 51k 3D object models in its curated form. Still, it does not offer enough breadth and detail to allow for training open-vocabulary reconstruction models and as a result, most proposed reconstruction methods focused on few specific categories, such as cars or chairs [choy_3d-r2n2_2016, wang_pixel2mesh_2018, rematas_sharf_2021]. More recently, OmniObjects3D [wu_omniobject3d_2023] and GSO [downs_google_2022] proposed highly-detailed 3D object datasets, albeit at a smaller scale than ShapeNet. In 2023, Objaverse-XL [deitke_objaverse-xl_2023] presented an inflection point for 3D datasets, offering over 10 million highly-diverse 3D models. For the first time, 3D reconstruction models trained on this dataset made the value proposition of being able to reconstruct ’any’ object [liu_zero-1--3_2023]. A recurring theme among large-scale 3D datasets is the reliance on synthetic object rendering. With robotics being an area that is heavily concerned with the Sim2Real gap, including subtle differences between renderings and real-world camera footage [jaunet_sim2realviz_2021], relying solely on simulated training data might not translate to good downstream performance.
Scene Datasets
@@
Regarding scene-level datasets, the focus is not on object diversity but rather on diversity of scene layout and composition. Similar to object-centric datasets, scene-level datasets can be synthetic and potentially procedurally generated [khanna_habitat_2024, kolve_ai2-thor_2022, deitke_procthor_2022, deitke_robothor_2020]. Procedural generation allows theoretically infinite scene layouts from a finite set of objects. Other datasets focus on manually obtained 3D scans of real scenes [chang_matterport3d_2017, baruch_arkitscenes_2021, sturm_benchmark_2012]. Due to the scanning process that is used for obtaining these datasets, one could say they are not truly 3D but only a collection of 2.5D information. As a result, the collected scenes frequently contain holes and do not feature complete object meshes. Instead, they only reconstruct what has been scanned. [xiang_posecnn_2018] provide a collection of 3D scenes featuring YCB objects [calli_ycb_2015], a collection of common household objects that are frequently featured in robot experiments [deng_self-supervised_2020, lu_systematic_2021, lerher_robotic_2023]. Recently, [pan_aria_2023] proposed a richly annotated digital twin dataset comprised of two indoor scenes.
III-G Applications in Robotics
In this section, we focus our attention on 3D reconstruction for creating digital twin environments to support robot manipulation. Reconstruction has had a major role in robotics research and a great many literature reviews have been published that discuss several areas in greater detail, including SLAM [yue_lidar-based_2024, al-tawil_review_2024], 3D Gaussian Splatting [zhu_3d_2024], as well as neural fields [slapak_neural_2024] for robotics tasks such as navigation and manipulation. We organise this section into two main categories. For completeness, we include dense- and multi-view reconstruction methods which have had a long history in robotics, while single-view reconstruction methods are more recent and the primary focus of the rest of this paper.
Dense- and Multi-View
@@ III-G Applications in Robotics
3D reconstruction from dense viewpoints has been extensively used in robotics, for example, SLAM is frequently used in mobile robotics and navigation [temeltas_slam_2008, cadena_past_2016, damjanovic_comprehensive_2025]. Closer to our real-to-sim setup, 3D Gaussians have recently been adopted for Real2Sim2Real in robotics [barcellona_dream_2024, han_re3sim_2025, li_robogsim_2024, lou_robo-gs_2024, wu_rl-gsbridge_2024, qureshi_splatsim_2024, zhu_vr-robo_2025, jia_discoverse_2024]. Typically, 3D Gaussians are learnt alongside coupled mesh representations of objects from dense view points. The 3D Gaussians are used for realistic rendering to close the sim-to-real gap, while the mesh-based representations are used for rigid body simulation. The resulting pipeline can be used for trajectory augmentation and reinforcement learning in simulation. Instead of a dual mesh representation, [abou-chakra_physically_2024] use a particle-based representation and simulator while also relying on 3D Gaussians for rendering fidelity.
@@
[torne_reconciling_2024] propose a similar pipeline but instead of relying on dual representations tied to 3D Gaussians, they only reconstruct scene meshes by densely scanning the scene and relying on human annotations. Again, the pipeline is used for reinforcement learning in simulation. Similarly, [patel_real--sim--real_2024] also forego Gaussian Splatting to instead reconstruct object meshes obtained from dense input viewpoints. [pfaff_scalable_2025] technically only require a single camera viewpoint for reconstructing mesh representations from object observations but rely on a robotic arm to present the object from different angles and re-grasp it to reduce occlusions.
@@
While certainly pushing the state-of-the-art of real-to-sim capabilities in robotics, these methods all require substantial scene coverage in their input data. This is a major limitation that as of yet has to be critically examined in the literature. Dense observations are not always available in real-world scenarios, especially when dealing with unstructured environments or when the robot is required to operate quickly. This limitation highlights the need for reconstruction methods that can effectively operate with limited input data, such as single-view reconstruction methods.
Single-View
@@
Recent open-vocabulary single-view reconstruction methods have been proposed for robotics applications, but the focus has been very limited. Existing works face several limitations, such as having to learn dynamics models, not reconstructing exact digital twins, or only considering the reconstructed scene as a static entity without simulation.
@@
[mu_robotwin_2024] propose generating digital twins from single image observations for robotic learning. However, they do not produce exact digital twins but instead insert a language and 2D image generation model into their pipeline. These are then used for generating diverse sets of objects similar to the observed objects, which are assembled into domain-randomised environments for use in downstream tasks. In addition to not producing exact digital twins, their method relies on a commercial reconstruction model, is only evaluated on very simple tabletop scenes without occlusion, and does not explicitly address object collisions resulting from reconstruction. [agarwal_scenecomplete_2024] pose the question whether single-view object-centric 3D reconstruction of entire scenes can be used in robotics. They show that using only pre-trained models, one can produce digital twins that afford computing parallel grasp poses that transfer to the real-world scene. However, they do not perform any form of simulation on the reconstructed scene. [yao_cast_2025] propose generating digital twins of entire scenes from single-view observation while also addressing physical constraints such as object collision and stability. Their process relies on GPT-4v [openai_gpt-4_2024] to produce accurate scene graphs that are used in their gradient-based physics optimisation. While stating that their approach enables Real2Sim transfer for robotics, the authors do not provide any experimental evaluation for this claim. Similarly, [katara_gen2sim_2024] mention their generative approach for domain randomisation could be used for creating digital twin environments for robotics but no further details or experimental evaluations are provided.
@@
Different from the above mesh reconstruction approaches, [lu_manigaussian_2025] produce 3D Gaussians while only requiring a single input view. However, as existing robotics simulators do not support 3D Gaussians, the authors propose to learn a neural dynamics model instead.
@@
Our work addresses an important gap in the literature by conducting a rigorous empirical evaluation of single-view reconstruction methods specifically in the context of robotics requirements. We focus on evaluating whether recent advances in computer vision for single-view 3D reconstruction can actually meet the particular needs of robotics simulation, providing concrete evidence of their current limitations and identifying specific areas for improvement.
@@
[chatterjee_3d-scene-former_2024] Single Pretrained Scene Mesh Feed-Forward Yes Object Collision No
[xu_agg_2024] Single No Object 3D Gaussians Feed-Forward No No No
[wu_amodal3r_2025] Single No Object, Scene 3D Gaussians, Radiance Field, Mesh Feed-Forward Segments single foreground object No No
[szymanowicz_bolt3d_2025] Single, Multi No Scene 3D Gaussians Feed-Forward No No No
[chu_buol_2023] Single Pretrained Scene Voxel Feed-Forward Yes Objects cannot collide as scene is represented by voxels Yes
[yao_cast_2025] Single Pretrained Scene Mesh Iterative Feed-Forward (Object Reconstruction), GPT-4v (Scene Graph), Gradient-Descent (Physics Optimisation) Yes Collision and Stability No
[zhang_clay_2024] Single Yes Object Mesh Feed-Forward No No No
[wu_direct_2024] Single Custom Object 3D Gaussians, Mesh Feed-Forward No No No
[xu_dmv3d_2023] Single No Object NeRF, Mesh Feed-Forward No No No
[sun_dreamcraft3d_2023] Single Pretrained Object Implicit Surface, Mesh Gradient-Descent (Score Distillation) No No Yes
[tang_dreamgaussian_2023] Single No Object 3D Gaussians, Mesh Gradient-Descent (Score Distillation & Texture Refinement) No No Yes
[li_dso_2025] Single No Object Mesh Feed-Forward No Fine-tuned to predict stable object Yes
[szymanowicz_flash3d_2024] Single Pretrained Scene 3D Gaussians Feed-Forward No No Yes
[shen_gamba_2024] Single No Object 3D Gaussians Feed-Forward No No Yes
[dogaru_generalizable_2024] Single Pretrained Scene Mesh Gradient-Descent (Score Distillation) by default; can be combined with other models Yes No Yes
[xu_grm_2025] Single No Object 3D Gaussians, Mesh Feed-Forward No No No
[zhang_holistic_2021] Single No Scene SDF, Mesh Feed-Forward, Iterative Refinement Yes Collision Yes
[melas-kyriazi_im-3d_2024] Single No Object 3D Gaussians Feed-Forward (Video Gen), Gradient-Descent (3D Gaussians) No No No
[li_instant3d_2023] Text, Single, Multi No Object Triplane, NeRF, Mesh Feed-Forward No No No
[xu_instantmesh_2024] Single No Object Mesh Feed-Forward No No Yes
[lu_large_2024] Single No Object 3D Gaussians Feed-Forward No No No
[tang_lgm_2025] Single No Object 3D Gaussians, Mesh Feed-Forward No No Yes
[hong_lrm_2023] Single No Object NeRF, Mesh Feed-Forward No No No
[qian_magic123_2023] Single No Object Mesh Gradient-Descent (Score Distillation) No No Yes
[wei_meshlrm_2025] Single No Object Mesh Feed-Forward No No No
[zhao_michelangelo_2023] Single No Object Mesh Feed-Forward No No Yes
[huang_midi_2024] Single No Scene SDF, Mesh Feed-Forward Yes Soft Collision Yes
[wu_multiview_2023] Single Yes Object, Scene Point Cloud Feed-Forward No No Yes
[shi_mvdream_2023] Single No Object NeRF Gradient-Descent (Score Distillation) No No Yes
@@
[wen_object-aware_2023] Single Pretrained Scene Mesh Feed-Forward Yes Collision No
@@
[liu_one-2-3-45_2023] Single No Object SDF, Mesh Feed-Forward No No Yes
[liu_one-2-3-45_2024] Single No Object SDF, Mesh Feed-Forward No No No
[yan_phycage_2024] Single No Object 3D Gaussians Feed-Forward (Multi-View Gen), Gradient-Descent (SDF, Score Distillation, Diff’ble Physics Simulation) Single object into 2 components Full Simulation Yes
[chen_physgen3d_2025] Single Pretrained Scene Mesh Feed-Forward (Reconstruction), GPT-4o, Gradient-Descent (Diff’ble Renderer) Yes Reconstructions can be used in particle simulator but stability and collision play no role during reconstruction process. Yes
@@
[mezghanni_physical_2022] Single No Object SDF Feed-Forward (SDF), Gradient-Descent (Diff’ble Physics Sim) No Stability No
[guo_physically_2024] Single No Object Mesh Gradient-Descent No Stability, Mechanical Material Properties Yes
[jiang_real3d_2024] Single No Object Mesh Feed-Forward No No Yes
[han_reparo_2024] Single Pretrained Scene Mesh Gradient-Descent (Score Distillation & Diff’ble Rendering for Poses) Yes No Not functional
[boss_sf3d_2024] Single No Object Mesh Feed-Forward No No Yes
[jun_shap-e_2023] Single No Object NeRF, Mesh Feed-Forward No No Yes
[chen_single-view_2024] Single No Scene Mesh Feed-Forward Yes No Yes
[szymanowicz_splatter_2024] Single No Object 3D Gaussians Feed-Forward No No Yes
[voleti_sv3d_2025] Single No Object Mesh Feed-Forward (Multi-View Gen), Gradient-Descent (Score Distillation) No No Yes
@@
[liu_syncdreamer_2023] Single No Object Mesh Gradient-Descent (Score Distillation) No No Yes
[chen_transformers_2022] Single No Object NeRF Feed-Forward No No Yes
[zou_triplane_2024] Single No Object 3D Gaussians Feed-Forward No No Yes
[wu_unique3d_2024] Single Custom Object Mesh Feed-Forward No No Yes
[qi_vpp_2023] Single No Object Point Cloud Feed-Forward No No Yes
[long_wonder3d_2024] Single No Object Mesh Feed-Forward (Multi-View Gen), Gradient-Descent (SDF) No No Yes
@@
[liu_zero-1--3_2023] Single No Object Mesh Gradient-Descent (Score Distillation) No No Yes
@@
[huang_zeroshape_2024] Single Custom Object Mesh Feed-Forward No No Yes

### v2

@@ III-E Physics
Physical plausibility is essential for physics simulation. For rigid bodies, this entails aspects of the overall scene composition, such as object collision and stability, as well as physical properties of individual objects themselves.
@@
On the scene level, many approaches formulate loss terms that are incorporated into gradient-based optimisation. For instance, SDF-based representations allow querying whether a particular point in space belongs to the volume of more than one object. As the SDF representation itself is typically optimised with gradient descent, adding a penalty term for doubly occupied space into the overall optimisation is straightforward [156, 157, 158]. [78] and [159] also regularise collision between objects and scene boundaries and background walls by relying on SDFs and bounding boxes, respectively. [160] and [161] take a slightly different approach by not discouraging existing collisions between objects, but instead clustering and decomposing a scene representation into individual objects which naturally addresses object overlap. [36] propose a complex setup that first queries a vision-language model to produce a scene graph containing object-object relations, such as one object supporting another or objects generally being in close contact with one another. Based on these pairwise graph relations, the authors select loss terms from a set of handcrafted loss functions that are then used to optimise physical scene coherence via gradient descent. [154] do not formulate explicit loss terms but instead resolve collisions by coupling the denoising process of multiple objects in their proposed 3D diffusion model, leading to fewer violations.
@@
Instead of handcrafted loss-terms, some approaches resort to differentiable physics simulation. This allows penalising anything that would result in undesirable behaviour during physics simulation, including object collision and object stability - even involving multiple objects at once. [75] dissect single objects into several components, such as a frog wearing a scarf, and optimise component shapes to resolve collision and object stability. [72] perform a similar process on the scene level by reconstructing individual objects and adapting their position and shape such that the resulting scene is physically stable.
@@
Some works do not consider entire scenes or compositional objects but instead concentrate on individual objects only. [94] learn a stability predictor that is used for guiding object reconstruction. [73] formulate differentiable loss terms to optimise stability of single objects. [74] use differentiable simulation during text-to-3D generation for generating stable objects, and [162] augment an SDF decoder with a differentiable simulator for object stability.
@@
Beyond stability and collision, researchers have also looked at estimating object articulations [163, 164, 165] and material properties [80, 166, 73]. While certainly useful for digital twinning of more complex environments, these aspects go beyond the scope of our study of whether reconstructing physically coherent scenes in simulation for robotics is feasible with current category-agnostic 3D reconstruction methods.
@@ III-F Training Data
3D datasets have historically been smaller than text or image datasets and most existing large-scale 3D datasets are primarily object-centric rather than scene-centric. This focus on isolated objects fails to capture important contextual relationships between objects in realistic environments. For robotics applications seeking to create accurate digital twins of entire scenes, having training data that represents realistic object arrangements, support relationships, and partial occlusions is crucial. Scene-level data would allow reconstruction models to learn typical object configurations and physical constraints that govern real-world environments. Without such contextual learning, models struggle to produce physically plausible reconstructions of complex scenes, leading to the stability and collision issues we observe in our evaluations.
@@
Object Datasets. ShapeNet [45] has been very influential ever since its introduction, offering around 51k 3D object models in its curated form. Still, it does not offer enough breadth and detail to allow for training category-agnostic reconstruction models and as a result, most proposed reconstruction methods focused on few specific categories, such as cars or chairs [115, 90, 136]. More recently, OmniObjects3D [47] and GSO [167] proposed highly-detailed 3D object datasets, albeit at a smaller scale than ShapeNet. In 2023, Objaverse-XL [48] presented an inflection point for 3D datasets, offering over 10 million highly-diverse 3D models. For the first time, 3D reconstruction models trained on this dataset made the value proposition of being able to reconstruct ’any’ object [70]. A recurring theme among large-scale 3D datasets is the reliance on synthetic object rendering. With robotics being an area that is heavily concerned with the Sim2Real gap, including subtle differences between renderings and real-world camera footage [168], relying solely on simulated training data might not translate to good downstream performance.
@@
Scene Datasets. Regarding scene-level datasets, the focus is not on object diversity but rather on diversity of scene layout and composition. Similar to object-centric datasets, scene-level datasets can be synthetic and potentially procedurally generated [169, 170, 171]. Procedural generation allows theoretically infinite scene layouts from a finite set of objects. Other datasets focus on manually obtained 3D scans of real scenes [172, 173]. Due to the scanning process that is used for obtaining these datasets, one could say they are not truly 3D but only a collection of 2.5D information. As a result, the collected scenes frequently contain holes and do not feature complete object meshes. Instead, they only reconstruct what has been scanned. [38] provide a collection of 3D scenes featuring YCB objects [174], a collection of common household objects that are frequently featured in robot experiments [175, 176, 177]. Recently, [39] proposed a richly annotated digital twin dataset comprised of two indoor scenes.
III-G Recent Uses of 3D Reconstruction in Robotic Systems
In this section, we review recent works that incorporate 3D reconstruction into robotic systems, with a particular focus on digital twin generation as a common use case. 3D reconstruction has played a central role in robotics research, serving as a foundation for tasks such as navigation, manipulation, and scene understanding. Numerous surveys have addressed specific technical areas in detail, including SLAM [178, 179], 3DGS [57], as well as neural fields [180]. While navigation and mobile robotics are commonly handled with SLAM, this survey focuses primarily on manipulation tasks, where digital twins are used for interaction and simulation. Rather than providing an exhaustive taxonomy of robotic tasks, we aim to examine how complete-geometry 3D reconstruction is being used to support downstream robotics capabilities, such as manipulation, grasp planning, and simulation-based policy learning.
3D reconstructions often act as a bridge between perception and control - used for building simulation environments, enabling collision checking, supporting object re-identification, or computing affordances. In this sense, we treat the integration of 3D reconstruction in robotic manipulation pipelines as the "application" of interest.
To structure our discussion, we group recent work into two categories: those using dense- or multi-view reconstruction, typically assuming extensive sensor coverage, and those using single-view reconstruction, which are more constrained and thus more aligned with real-time robotic settings.
@@ III-G Recent Uses of 3D Reconstruction in Robotic Systems
Dense- and Multi-View. 3D Gaussians have recently been adopted for Real2Sim2Real in robotic manipulation [15, 16, 17, 18, 19, 20, 22]. Typically, 3D Gaussians are learnt alongside coupled mesh representations of objects from dense view points. The 3D Gaussians are used for realistic rendering to close the sim-to-real gap, while the mesh-based representations are used for rigid body simulation. The resulting pipeline can be used for trajectory augmentation and reinforcement learning in simulation. Instead of a dual mesh representation, [181] use a particle-based representation and simulator while also relying on 3D Gaussians for rendering fidelity.
@@
[23] propose a similar pipeline but instead of relying on dual representations tied to 3D Gaussians, they only reconstruct scene meshes by densely scanning the scene and relying on human annotations. Again, the pipeline is used for manipulation through reinforcement learning in simulation. Similarly, [24] also forego 3DGS to instead reconstruct object meshes obtained from dense input viewpoints. [25] technically only require a single camera viewpoint for reconstructing mesh representations from object observations but rely on a robotic arm to present the object from different angles and re-grasp it to reduce occlusions.
@@
While certainly pushing the state-of-the-art of real-to-sim capabilities in robotic manipulation, these methods all require substantial scene coverage in their input data. This is a major limitation that as of yet has to be critically examined in the literature. Dense observations are not always available in real-world scenarios, especially when dealing with unstructured environments or when the robot is required to operate quickly. This limitation highlights the need for reconstruction methods that can effectively operate with limited input data, such as single-view reconstruction methods.
@@
Single-View. Recent category-agnostic single-view reconstruction methods have been proposed for robotic manipulation, but the focus has been very limited. Existing works face several limitations, such as having to learn dynamics models, not reconstructing exact digital twins, or only considering the reconstructed scene as a static entity without simulation.
@@
[34] propose generating digital twins from single image observations for robotic manipulation. However, they do not produce exact digital twins but instead insert a language and 2D image generation model into their pipeline. These are then used for generating diverse sets of objects similar to the observed objects, which are assembled into domain-randomised environments for use in downstream tasks. In addition to not producing exact digital twins, their method relies on a commercial reconstruction model, is only evaluated on very simple tabletop scenes without occlusion, and does not explicitly address object collisions resulting from reconstruction. [35] pose the question whether single-view object-centric 3D reconstruction of entire scenes can be used in robotic manipulation. They show that using only pre-trained models, one can produce digital twins that afford computing parallel grasp poses that transfer to the real-world scene. However, they do not perform any form of simulation on the reconstructed scene. [36] propose generating digital twins of entire scenes from single-view observation while also addressing physical constraints such as object collision and stability. Their process relies on GPT-4v [182] to produce accurate scene graphs that are used in their gradient-based physics optimisation. While stating that their approach enables Real2Sim transfer for robotics, the authors do not provide any experimental evaluation for this claim. Similarly, [37] mention their generative approach for domain randomisation could be used for creating digital twin environments for robotics but no further details or experimental evaluations are provided.
@@
Different from the above mesh reconstruction approaches, [183] produce 3D Gaussians while only requiring a single input view. However, as existing robotics simulators do not support 3D Gaussians, the authors propose to learn a neural dynamics model instead.
@@
Our work addresses an important gap in the literature by conducting a rigorous empirical evaluation of single-view reconstruction methods specifically in the context of robotic manipulation requirements. We focus on evaluating whether recent advances in computer vision for single-view 3D reconstruction can actually meet the particular needs of robotics simulation for manipulation, including accuracy and physical plausibility, and provide concrete evidence of their current limitations and identifying specific areas for improvement.
@@
[159] Single Pretrained Scene Mesh Feed-Forward Yes Object Collision No
[101] Single No Object 3D Gaussians Feed-Forward No No No
@@
[184] Single No Object, Scene 3D Gaussians, Radiance Field, Mesh Feed-Forward Segments single foreground object No No
[152] Single, Multi No Scene 3D Gaussians Feed-Forward No No No
[185] Single Pretrained Scene Voxel Feed-Forward Yes Objects cannot collide as scene is represented by voxels Yes
@@
[36] Single Pretrained Scene Mesh Iterative Feed-Forward (Object Reconstruction), GPT-4v (Scene Graph), Gradient-Descent (Physics Optimisation) Yes Collision and Stability No
[186] Single Yes Object Mesh Feed-Forward No No No
[187] Single Custom Object 3D Gaussians, Mesh Feed-Forward No No No
[143] Single No Object NeRF, Mesh Feed-Forward No No No
@@
[188] Single Pretrained Object Implicit Surface, Mesh Gradient-Descent (Score Distillation) No No Yes
@@
[189] Single No Object 3D Gaussians, Mesh Gradient-Descent (Score Distillation & Texture Refinement) No No Yes
@@
[190] Single No Object Mesh Feed-Forward No Fine-tuned to predict stable object Yes
[102] Single Pretrained Scene 3D Gaussians Feed-Forward No No Yes
[191] Single No Object 3D Gaussians Feed-Forward No No Yes
[153] Single Pretrained Scene Mesh Gradient-Descent (Score Distillation) by default; can be combined with other models Yes No Yes
[100] Single No Object 3D Gaussians, Mesh Feed-Forward No No No
[157] Single No Scene SDF, Mesh Feed-Forward, Iterative Refinement Yes Collision Yes
[98] Single No Object 3D Gaussians Feed-Forward (Video Gen), Gradient-Descent (3D Gaussians) No No No
[123] Text, Single, Multi No Object Triplane, NeRF, Mesh Feed-Forward No No No
[27] Single No Object Mesh Feed-Forward No No Yes
[192] Single No Object 3D Gaussians Feed-Forward No No No
[96] Single No Object 3D Gaussians, Mesh Feed-Forward No No Yes
[135] Single No Object NeRF, Mesh Feed-Forward No No No
[79] Single No Object Mesh Gradient-Descent (Score Distillation) No No Yes
[193] Single No Object Mesh Feed-Forward No No No
[28] Single No Object Mesh Feed-Forward No No Yes
[154] Single No Scene SDF, Mesh Feed-Forward Yes Soft Collision Yes
[194] Single Yes Object, Scene Point Cloud Feed-Forward No No Yes
[127] Single No Object NeRF Gradient-Descent (Score Distillation) No No Yes
[195] Single Pretrained Scene Mesh Feed-Forward Yes Collision No
[26] Single No Object SDF, Mesh Feed-Forward No No Yes
[81] Single No Object SDF, Mesh Feed-Forward No No No
[75] Single No Object 3D Gaussians Feed-Forward (Multi-View Gen), Gradient-Descent (SDF, Score Distillation, Diff’ble Physics Simulation) Single object into 2 components Full Simulation Yes
[196] Single Pretrained Scene Mesh Feed-Forward (Reconstruction), GPT-4o, Gradient-Descent (Diff’ble Renderer) Yes Reconstructions can be used in particle simulator but stability and collision play no role during reconstruction process. Yes
[162] Single No Object SDF Feed-Forward (SDF), Gradient-Descent (Diff’ble Physics Sim) No Stability No
[73] Single No Object Mesh Gradient-Descent No Stability, Mechanical Material Properties Yes
[71] Single No Object Mesh Feed-Forward No No Yes
[155] Single Pretrained Scene Mesh Gradient-Descent (Score Distillation & Diff’ble Rendering for Poses) Yes No Not functional
[80] Single No Object Mesh Feed-Forward No No Yes
[134] Single No Object NeRF, Mesh Feed-Forward No No Yes
[197] Single No Scene Mesh Feed-Forward Yes No Yes
[99] Single No Object 3D Gaussians Feed-Forward No No Yes
[198] Single No Object Mesh Feed-Forward (Multi-View Gen), Gradient-Descent (Score Distillation) No No Yes
[69] Single No Object Mesh Gradient-Descent (Score Distillation) No No Yes
[122] Single No Object NeRF Feed-Forward No No Yes
[145] Single No Object 3D Gaussians Feed-Forward No No Yes
[129] Single Custom Object Mesh Feed-Forward No No Yes
[199] Single No Object Point Cloud Feed-Forward No No Yes
[200] Single No Object Mesh Feed-Forward (Multi-View Gen), Gradient-Descent (SDF) No No Yes
[70] Single No Object Mesh Gradient-Descent (Score Distillation) No No Yes
[142] Single Custom Object Mesh Feed-Forward No No Yes